OpenStudy (anonymous):

The equation p=2.50v-500 gives the estimated profit "p" in dollars if v people visit the park In parts (a)-(b) find the answer by writing and solving an equation or inequality. a.For what number of visitors will the profit be about $2ooo b. 200 people visit. What is about the concession-stand profit for the day? c.For what number of visitors will the profit be at least $500 PLEASE HELP AND SHOW WORK :)

OpenStudy (anonymous):

a) replace P by 2000 solve for v in two steps add 500, divide by 2.5

OpenStudy (anonymous):

b) replace v by 200 and compute P

OpenStudy (anonymous):

c) solve \[2.5v-500\geq 500\] by adding 500 to both sides and then dividing by 2.5

OpenStudy (callisto):

a) put P=2000 into the equation 2000 = 2.50v - 500 2500 = 2.50v v= 1000 b) Put v = 200 into equation p= 2.50 x 200 - 500 = 0 c) like what satellite73 said. Solve it and you'll get \[v \ge400\]